OpenStudy (anonymous):

Factor this polynomial completely over the set of complex numbers. x^3-x^2-4x+4

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

by inspection of the coefficients, 1 is a zero, so x - 1 is a factor x^3 - x^2 - 4x + 4 = (x - 1)(x^2 - 4) you can factor by grouping: x^2 (x - 1) - 4(x - 1) = (x - 1)(x^2 - 4) and x^2 - 4 = (x - 2)(x + 2) fully factored: (x - 1)(x - 2)(x + 2) zeros: 1 , 2 , -2 all three zeros are real (or complex with an imaginary component of 0)